OpenStudy (anonymous):

write a system of equations to model the situation. solve by any method. Sarah is 25 years older than her son gavin. In ten years, SArah willb e twice Gavin's age. How old are sarah and gavin now?

OpenStudy (anonymous):

s = g + 25 (s + 10) = 2(g + 10) Those are the simultaneous equations. Can you solve them from here?

OpenStudy (anonymous):

yes,i believe i can try! thank you very much!

OpenStudy (anonymous):

If you substitute the first equation into the second: (g + 25) + 10 = 2(g + 10) g + 35 = 2g + 20 g = 15 so s = 40

OpenStudy (anonymous):

You're welcome! Good luck!

OpenStudy (anonymous):

Good luck and thx for the recognition!

OpenStudy (anonymous):

thank you and you deserve it!